OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tangent of seven pi divided by eight.

OpenStudy (anonymous):

\[\sin(x)=\sqrt{\frac{ 1-\cos(2x) }{ 2 }}\]\[\cos(x)=\sqrt{\frac{ 1+\cos(2x) }{ 2}}\]then\[\tan(x)=\sqrt{\frac{ 1-\cos(2x) }{ 1+\cos(2x) }}\] and\[\cos(2x)=\cos(7\pi/4)=\cos(2 \pi-\pi/4)=\cos(-\pi/4)=\cos(\pi/4)=\frac{ \sqrt{2} }{ 2 }\]Then you have:\[\tan(7\pi/8)=\sqrt{\frac{ 2-\sqrt{2} }{ 2+\sqrt{2}}}=0.4142\]Now pay attention to the sign as 7*pi/8 is in the second quadrant, then its tangent has to be negative. Then your solution is:-0.4142

OpenStudy (anonymous):

oh wow so the answer is -0.4142?

OpenStudy (anonymous):

exactly, check it with your calculator

OpenStudy (anonymous):

kk wil do one sec I did that horribly wrond then -_-

OpenStudy (anonymous):

sure as poopits right thank you so much

OpenStudy (anonymous):

u r w